in right angle triangle pqr are pq=7cm,qr=25cm and angle q=90 find tan q-tanr

Grade:10

1 Answers

Vikas TU
7 years ago
I think angle is P that is 90 degree.
q = 90 degree will lead to contradiction according to the question.
So assuming P = 90 degree,
 
 
PQ = 7cm and QR = 25 cm.
PQR is a right angle triangleBy using Pythagoras theorem
PR2 =QR2 - PQ2
PR2=576
PR=24
Tan Q=PR/PQ=24/7
Tan R=PQ/PR=7/24
Tan Q-Tan R=527/168